### v1

# Simulation to Rules: A Dual-VLM Framework for Formal Visual Planning

Abstract
Vision Language Models (VLMs) show strong potential for visual planning but struggle with precise spatial and long-horizon reasoning. In contrast, Planning Domain Definition Language (PDDL) planners excel at long-horizon formal planning, but cannot interpret visual inputs. Recent works combine these complementary advantages by enabling VLMs to turn visual planning problems into PDDL files for formal planning. However, while VLMs can generate PDDL problem files satisfactorily, they struggle to accurately generate the PDDL domain files, which describe all the planning rules. As a result, prior methods rely on human experts to predefine domain files or on constant environment access for refinement. We propose VLMFP, a Dual-VLM-guided framework that can autonomously generate both PDDL problem and domain files for formal visual planning. VLMFP introduces two VLMs to ensure reliable PDDL file generation: A SimVLM that simulates action consequences based on input rule descriptions, and a GenVLM that generates and iteratively refines PDDL files by comparing the PDDL and SimVLM execution results. VLMFP unleashes multiple levels of generalizability: The same generated PDDL domain file works for all the different instances under the same problem, and VLMs generalize to different problems with varied appearances and rules. We evaluate VLMFP with 6 grid-world domains and test its generalization to unseen instances, appearance, and game rules. On average, SimVLM accurately describes 95.5%, 82.6% of scenarios, simulates 85.5%, 87.8% of action sequence, and judges 82.4%, 85.6% goal reaching for seen and unseen appearances, respectively. With the guidance of SimVLM, VLMFP can generate PDDL files to reach 70.0%, 54.1% valid plans for unseen instances in seen and unseen appearances, respectively. Project page: https://sites.google.com/view/vlmfp.
1 Introduction
Although Large Language Models (LLMs) have shown strong performance in solving text-based planning problems (Wei et al., 2022; Yao et al., 2022; Raman et al., 2022; Yao et al., 2024), many real-world planning tasks, such as robot assembly, drone navigation, and autonomous driving, are inherently visual, making the reliance on carefully engineered text inputs impractical and limiting. This gap motivates the shift toward VLM-based planning, where visual inputs provide a more direct and intuitive basis for reasoning. However, current VLMs lack precise spatial understanding and long-horizon reasoning, which constrains their ability to address complex, multi-step planning problems that involve intricate spatial relationships among multiple objects (Wu et al., 2024).
On the other hand, Planning Domain Definition Language (PDDL) (McDermott, 2000) is a formal language designed to describe planning problems and domains in a structured, machine-interpretable way. PDDL has enabled numerous automated planners to derive long-horizon solutions. However, although PDDL-based planners excel at reasoning over structured domains, they depend on correctly structured PDDL domain and problem files and cannot directly interpret visual inputs. Constructing accurate PDDL definitions is non-trivial and requires expert knowledge, which is often inaccessible to non-expert users, limiting the broader adoption of PDDL planners in real-world scenarios.
Figure 1: An overview of VLMFP. Sections in orange are the inputs. Sections in green represents SimVLM and its outputs. Sections in blue denotes the GenVLM and its outputs. Sections in gray are provided. VLMFP takes in the image and domain description, describe the scenario, generate and validate PDDL files, compare their execution results with SimVLM, and update the PDDL files.
Recent works have explored combining the advantages of language models and PDDL planners through various approaches. (Liu et al., 2023; Xie et al., 2023) employ LLMs as translators to convert natural language scenario descriptions into PDDL problem files. (Mahdavi et al., 2024) leverages environment interactions to enable file generation for both problem and domain. However, these methods require either textual scenario descriptions, environment access, or pre-defined PDDL files, which cannot be directly achieved through visual inputs. More recently, vision-language models (VLMs) have been used to extract scenario information from visual inputs and generate corresponding PDDL problem files(Shirai et al., 2024; Dang et al., 2025). However, since current VLMs lack the ability to accurately generate the domain PDDL, these approaches also assume access to ground truth domain PDDL files, without which PDDL planners cannot produce any results.
In this paper, we address the challenge of visual long-horizon planning by introducing a novel framework, VLM-Guided Formal Planning (VLMFP, illustrated in Fig. 1), a Dual-VLM-guided framework that autonomously generates both problem and domain PDDL files for visual planning. VLMFP integrates two specialized VLMs: a fine-tuned SimVLM that perceives the scenario from visual inputs and simulates action outcomes, and a large GenVLM that generates and iteratively refines PDDL files by aligning their execution with SimVLM’s simulations. Generating both problem and domain PDDL from visual inputs requires object recognition, spatial understanding, reasoning, and PDDL knowledge. We fine-tune a small VLM as SimVLM to strengthen its spatial reasoning, while using a large model as GenVLM for general reasoning and extensive PDDL knowledge.
Importantly, VLMFP achieves multiple levels of generalizability. A generated domain PDDL can be reused across all instances of the same domain, while problem PDDL files can be adapted efficiently for new instances as in-context examples. The framework also transfers well to unseen appearances and even altered environment rules. We evaluate VLMFP on six grid world domains, showing that SimVLM reliably describes scenarios, simulates actions, and determines goal achievement, while GenVLM, guided by SimVLM feedback, generates valid PDDL files that enable planners to solve both seen and unseen problems.
In summary, our key contributions are:
  • We construct a large-scale dataset of 430k action sequence simulations with reasoning and feedback across 6 grid-world domains of different map sizes, appearances, and game rules. We fine-tune Qwen2-VL-7B with the dataset, and our finetuned model demonstrates strong generalization to unseen instances, appearances, and game rules.
  • We propose VLMFP, a Dual-VLM-guided framework that autonomously generates PDDL domain and problem files for visual planning, which, to our knowledge, is the first framework to leverage visual inputs to generate both PDDL files without human feedback or direct environment access.
[…]
2 Related Works
LLM and VLM Planning Although Large Language Models (LLMs) have achieved impressive performance on text-based planning tasks (Wei et al., 2022; Yao et al., 2022; Raman et al., 2022; Yao et al., 2024), many real-world planning problems are inherently visual. Relying solely on carefully constructed text inputs is often impractical and cannot capture the rich spatial details in visual environments. This limitation has motivated a growing interest in VLM-based planning, where visual observations provide a more direct and intuitive foundation for reasoning (Driess et al., 2023; Huang et al., 2023; Zhang et al., 2023; Nasiriany et al., 2024). However, current VLMs still fall short in precise spatial understanding (Wu et al., 2024), which restricts their effectiveness on complex planning tasks involving multiple objects and intricate spatial relationships. Some recent works explore Visual Chain-of-Thought (Li et al., 2025; Zhao et al., 2025), which use images as part of the reasoning process while generating the plan. However, this process introduces computational overhead and is hard to apply to long-horizon planning tasks. In this work, we fine-tune SimVLM to enhance its visual–spatial understanding and reasoning, enabling it to guide the generation of PDDL files for solving long-horizon planning problems.
LLM and VLM + PDDL Since existing LLMs lack the ability to reliably perform long-horizon reasoning in complex tasks (Achiam et al., 2023a; Valmeekam et al., 2022; 2023; Kambhampati et al., 2024), recent works have explored combining LLMs with external solvers to augment their reasoning and planning capabilities (Wu et al., 2022; He-Yueya et al., 2023; Pan et al., 2023; Ye et al., 2024; Li et al., 2023; 2024; Hao et al., 2024a; b). Among such solvers, combining LLMs with PDDL-based planners is a powerful option (Silver et al., 2022; Liu et al., 2023; Stein & Koller, 2023; Xie et al., 2023; Stein & Koller, 2023; Guan et al., 2023; Oswald et al., 2024; Mahdavi et al., 2024), as PDDL is designed for precise symbolic reasoning over long-horizon problems. However, these works either require predefined PDDL domain files, human corrections, or environment access to provide feedback while generating PDDL files. In addition, to further enable visual interpretation, many works combine VLMs with PDDL planners (Shirai et al., 2024; Dang et al., 2025) by generating PDDL Problem files based on image observation. However, similarly, as generating PDDL domain files is complicated, these works assume access to predefined domain files. In our work, we use a Dual-VLM-guided framework to autonomously generate both PDDL problem and domain files, without human guidance or environment access.
3 VLMFP
Figure 2: An example of how VLMFP tackle a Frozenlake instance.
3.1 Problem Formulation and Challenges
In our setting, a visual planning problem is defined by two pieces of information: ❶ A domain description $n_{d}$, which is a natural language description of the general rules, domain settings, available actions, and the planning objective of the problem; and ❷ a problem setup image $i_{p}$, which is an image showing the layouts, initial states of the problem instance.
Figure 2 top left shows an example domain description and problem setup image of the planning problem FrozenLake. As can be observed, the domain description elaborates the rules of the game (e.g., no stepping onto an ice hole) and available actions of the player (i.e., moving up, down, left, and right). The problem setup image is a map showing the initial positions of the player, the destination, and frozen likes.
The goal of VLMFP is, given the problem definition pair $(n_{d},i_{p})$, to come up with an action sequence to achieve the goal. Considering the complementary advantages of VLMs in image perception and natural language understanding, and PPDL in formal planning, VLMFP follows the two-step pipeline, where VLMs are adopted to translate problem definition into PDDL files, and then a PDDL planner is invoked to solve the problem.
However, an unresolved challenge of such a pipeline is that the translation into PDDL files using VLMs is an error-prone process. Specifically, PDDL provides a standardized framework for representing automated planning problems, which involves two files: ❶ A PDDL domain file $f_{d}$, which formally encodes the rules and domain descriptions by defining a set of predicates, actions with preconditions and effects; and ❷ a PDDL problem file $f_{p}$, which formally specifies objects, initial state, and goal conditions for a concrete problem instance. $f_{d}$ and $f_{p}$ can be roughly understood as the formal translation of the domain description $n_{d}$ and problem setup image $i_{p}$, respectively. Please refer to Appendix A.7 for example PDDL files.
The translations of these two files using VLMs come with their respective challenges. On one hand, translating the problem setup from the image $i_{p}$ to the PDDL problem file $f_{p}$ requires a precise understanding of the spatial relationship in the image, but existing VLMs tend to make mistakes in spatial perception tasks such as counting rows/columns, measuring distance etc (Wu et al., 2024). On the other hand, the PDDL domain file $f_{d}$ should be general enough to work for all problem instances under the same task (e.g., different maps of the same FrozenLake game), but generating the PDDL domain file using VLM alone, without human input or constant access to the environment, can easily fail to accurately model all constraints and dynamics(e.g., to move left, the goal cell must not be ice hole, and must be at left of the origin cell).
3.2 VLMFP Overview
To tackle the aforementioned challenges, VLMFP introduces a dual-VLM framework, consisting of a SimVLM (Simulation VLM) and a GenVLM (Generation VLM). SimVLM takes the domain description $n_{d}$, the problem setup image $i_{p}$, and a list of action $\pi=[a_{1:T}]$ (e.g., move left in the FrozenLake game), and performs three tasks: ❶ Describing the spatial relationship in the $i_{p}$ in natural language, denoted as $n_{p}$, ❷ reasoning and predicting the consequence of the proposal actions (e.g. whether the player hits a wall/ice hole), and ❸ determining whether executing $\pi$ successfully achieves the problem goal. GenVLM generates both PDDL files with the help of SimVLM.
The two VLMs should have different comparative advantages. SimVLM should be stronger at precisely understanding the spatial relationship in images, and have a superior capability in simulating action consequences. GenVLM should possess stronger general reasoning and question-answering capabilities, and have richer knowledge in PDDL.
With these comparative advantages, SimVLM is designed to support GenVLM in two ways. First, SimVLM’s generated description of the spatial relationships in the problem setup image $i_{p}$ is fed to GenVLM, therefore compensating the latter’s weakness in understanding spatial relationships. Second, and more importantly, SimVLM can verify the correctness of the PDDL files by comparing simulation results. Specifically, in VLMFP, there are two ways to predict the consequence of an action. The first is directly by SimVLM; the second is by execution based on the generated PDDL files. If the two results disagree, it indicates a greater probability of incorrect PDDL files, so the disagreement is provided to the GenVLM as feedback to refine the PDDL file generation.
Based on these rationales, VLMFP generates PDDL files via the following four steps:
Step 1: Candidate PDDL files generation. Given the domain descriptions and problem setup image, $(n_{d},i_{p})$, the SimVLM first generates a description of the spatial relationships ($n_{p}$) in the problem setup image, which is then fed to the GenVLM to generate candidate PDDL files, namely
$n_{p}=V_{S}(n_{d},i_{p}),\quad f^{(0)}_{d},f^{(0)}_{p}=V_{G}(n_{d},i_{p},n_{p}),$ (1)
where $V_{S}$ and $V_{G}$ represent the forward passes of SimVLM and GenVLM, respectively.
Step 2: Prescreening. The generated PDDL files are checked against syntactic correctness and semantic consistency.
Step 3: Simulation consistency checking. Random action sequences are executed in the PDDL environment based on the generated PDDL files, whose results are compared against the simulation results by SimVLM. Inconsistencies are summarized as feedback, denoted as $s$.
Step 4: PDDL files updating. GenVLM refines the generated PDDL files based on the feedback:
$f_{d}^{(t)},f_{p}^{(t)}=V_{G}\big(n_{d},i_{p},n_{p};s,f_{d}^{(t-1)},f_{p}^{(t-1)}\big).$ (2)
VLMFP iterates over steps 2-4 until consistency is achieved and a plan is found by the PDDL planner. Figure 2 gives an example of this process. In the following, Section 3.3 will discuss the implementation details of the two VLMs; Section 3.4 will provide further details of the above steps.
3.3 VLM Implementation Details
SimVLM. As discussed in Section 3.2, SimVLM is tasked with precisely describing the problem setup image $i_{p}$ and predicting action consequences. Since existing VLMs are generally weak in spatial relationship reasoning, we fine-tune Qwen2-VL-7B to accomplish these tasks. Specifically, we collected six grid world domains of varying complexity. For each domain, we collect data across map sizes ranging from 3 to 8 with varying obstacle probabilities, and create 5–6 distinct visual appearances. In total, this yields a dataset of 430k datapoints.
The fine-tuning process enables SimVLM to better align visual observations with spatially grounded reasoning. In particular, SimVLM learns to generate concise natural language narratives of the initial scenario and to simulate the outcomes of action sequences within these domains. By grounding visual inputs in structured textual descriptions and action simulations, SimVLM serves as a reliable intermediate model that bridges raw perception and the formal representation required for PDDL generation. We show examples of open-source large VLMs’ failure to accurately describe the scenario and simulate actions in Appendix A.4.1.
GenVLM. While SimVLM is tailored for perception and structured simulation, the generation of PDDL domain and problem files requires advanced reasoning to capture action dynamics and a precise understanding of PDDL syntax and conventions to ensure valid domain specifications. To address this, we leverage a large VLM API, GPT-4o (Achiam et al., 2023b), referred to as GenVLM, which has broader reasoning capacity and linguistic knowledge needed for reliable PDDL construction. Guided by the scenario descriptions and simulated outcomes provided by SimVLM, GenVLM generates initial PDDL files and iteratively refines them to resolve inconsistencies. In this way, GenVLM complements SimVLM by bridging high-level reasoning with formal symbolic representation.
3.4 Algorithmic Details
Prescreening. In this step, VLMFP generates and verifies whether the generated initial PDDL problem and domain files are structurally valid before moving on to further evaluation. A PDDL domain or problem is valid if it is syntactically correct and semantically consistent, meaning the domain’s actions, predicates, and types are well-defined and the problem’s objects, initial state, and goals align with that domain.
For example, as shown in Fig. 2, in the FrozenLake domain, the problem PDDL defines ice hole positions using predicate ice-hole, but if the Domain PDDL does not define ice-hole under the (:predicates section, these two files would be identified as invalid and cannot pass the prescreening. We set the maximum rounds of file regeneration to be 5. This stage is crucial because it filters out syntactic or structural errors at an early step, thereby ensuring that only valid PDDL files progress to the consistency checking process and that subsequent comparisons are meaningful.
Simulation consistency checking. In this step, VLMFP evaluates the fidelity of the generated PDDL files by comparing their execution results with those of SimVLM. Random executable action sequences are sampled in both environments and executed in the other environment: SimVLM simulates the outcomes and provides reasoning step by step, and the PDDL environment checks action executability under the generated domain.
For a walk length $T$, let $P_{\mathrm{sim},T}$ be a distribution over SimVLM-executable $T$-step action sequences, where $E_{f_{d},f_{p}}(q)\in\{0,1\}$ indicates whether $q$ is executable in the generated PDDL environment $(f_{d},f_{p})$. Similarly, let $P_{f_{d},f_{p},T}$ be a distribution over $T$-step walks executable in $(f_{d},f_{p})$, and let $E_{\mathrm{sim}}(q)\in\{0,1\}$ indicate SimVLM executability. Following (Mahdavi et al., 2024), we define the Exploration Walk (EW) score as:
$m_{\mathrm{EW}}(\hat{d},\hat{p})=2\bigg(\Big({\frac{1}{T_{\max}}\sum_{T=1}^{T_{\max}}\mathbb{E}_{q\sim P_{\mathrm{sim},T}}[E_{f_{d},f_{p}}(q)]}\Big)^{-1}+\Big({\frac{1}{T_{\max}}\sum_{T=1}^{T_{\max}}\mathbb{E}_{q\sim P_{f_{d},f_{p},T}}[E_{\mathrm{sim}}(q)]}\Big)^{-1}\bigg)^{-1}$ (3)
The EW score compares bi-directional similarity. A higher EW score indicates stronger alignment between the generated and reference environments. For example, in the FrozenLake domain, SimVLM predicts that moving right from pos-1-1 to pos-1-2 should fail due to an ice hole, whereas the generated PDDL environment instead allows the move. This discrepancy lowers the EW score and signals GenVLM to refine the domain file. Importantly, this stage not only enables VLMFP to detect execution mismatches and rate generated files using the EW score, but also produces natural language feedback on the incorrect actions to guide further file updates.
PDDL files updating. In this step, VLMFP refines the generated PDDL files based on the discrepancies identified in the last step. Using the feedback that describes the incorrect action and its expected outcome, GenVLM systematically inspects all critical components by listing and reasoning whether the objects, object types, init states, goal states, predicates, actions are valid and coherent, identifies which file(s) are erroneous, and applies targeted modifications, such as adding missing objects or correcting action preconditions, to regenerate updated PDDL specifications.
4 Experimental Results
4.1 Domains
Figure 3: Visualizations of some seen and an unseen appearances for six grid world domains.
We finetune SimVLM and test on six grid world domains: Frozenlake, Maze, Sokoban, Package, Printer, and Overcooked (Towers et al., 2024; Silver & Chitnis, 2020; Jin et al., 2023; Wu et al., 2021). For each domain, we collect data across map sizes ranging from 3 to 8 with varying obstacle probabilities, and create 5–6 distinct visual appearances. In total, this yields a dataset of 430k datapoints. By default, we use this dataset for SimVLM fine-tuning, which is used to evaluate framework VLMFP. Task descriptions, visualizations of different scenario appearances, and dataset statistics are included in Appendix A.2,  A.3, and  A.4.3.
4.2 SimVLM Performance and Appearance Generalization
We fine-tune a Qwen2-VL-7B model as SimVLM on datasets from six grid world domains. Given a domain description, an image, and an action sequence, SimVLM outputs (1) a natural language problem description, (2) step-by-step reasoning and outcomes, and (3) a judgment on goal achievement. Appendix A.4 provides fine-tuning details and sample inputs and outputs. Before integrating SimVLM into VLMFP, we evaluate its performance on these tasks individually. Since each datapoint has a fixed-format output, we measure performance by exact string matching between the model’s predictions and the ground-truth outputs. The four outputs are referred to as Task Description, Execution Reason, Execution Result, and Goal Reach.
To assess SimVLM’s generalization ability, we consider two settings: Seen (S) and Unseen (U) appearance. S evaluates images with visual styles present in training, while U evaluates novel appearances under the same rules. This measures how well SimVLM fits the training distribution and how robustly it transfers to new visual variations of the same task. For each domain, we evaluate on 1000 random datapoints for both S and U, and report the average string matching rate.
Table 1: String matching rate (%) for 4 SimVLM output types on 6 grid world domains.
Output Frozenlake Maze Sokoban Package Printer Overcooked Average
S U S U S U S U S U S U S U
Task Descrip. 100 92.3 99.1 89.8 74.9 51.6 100 99.7 99.9 96.3 99.2 65.6 95.5 82.6
Exec. Reason 96.3 96.0 85.3 97.1 76.6 84.6 89.8 88.4 87.4 86.8 78.9 75.9 85.7 88.1
Exec. Result 96.2 95.9 85.3 96.8 75.2 83.4 89.8 88.4 87.4 86.8 78.8 75.7 85.5 87.8
Goal Reach 94.0 93.7 80.3 94.7 74.2 83.2 87.8 86.0 84.1 84.6 73.9 71.2 82.4 85.6
Average 96.6 94.5 87.5 94.6 75.2 75.7 91.9 90.6 89.7 88.6 82.7 72.1 87.3 86.0
Results and Analysis. We summarize SimVLM’s performance on six grid-world domains under both seen (S) and unseen (U) appearance settings in Table 1. There are three key takeaways:
First, SimVLM achieves strong accuracy across tasks on seen visual appearances. On average, it reaches 95.5%, 85.7%, 85.5%, 82.4% in the seen setting across all domains for the four outputs, respectively. This highlights the effectiveness and reliability of SimVLM.
Second, SimVLM generalizes well to unseen visual appearances. On average, it reaches 82.6%, 88.1%, 87.8%, 85.6% in the unseen setting across all domains for the four outputs, respectively. The gap between S and U is minimal (average drop of 1.3% across all domains and outputs), showing that the fine-tuned model is not overfitting to training appearances and can transfer to new visual styles while maintaining stable output quality. This robustness is crucial for scaling to real-world environments where visual variation is common.
Third, errors in the Task Description do not necessarily affect action simulation. While the Task Description output under U drops by 12.9% on average compared to S, the execution simulation outputs remain equally accurate. This suggests that failures in Task Description typically involve only a small subset of objects, which often do not impact the sampled action sequences. For instance, the Sokoban task contains multiple types of objects: box, goal, wall, agent, floor, and the appearances of these objects are all changed. When the appearances of the objects in a domain change, since SimVLM is not familiar with the new look of objects, it could fail to recognize or localize some objects. However, when the misidentified objects are not directly involved in the sampled action sequence, execution reasoning and results remain correct. This separation highlights that SimVLM is robust in reasoning about action dynamics, even when its initial object recognition is imperfect.
Together, these results validate that SimVLM can produce reliable structured outputs from visual inputs and maintain strong generalization to new appearances, making it a suitable foundation for guiding PDDL generation in VLMFP.
4.3 VLMFP Performance
We evaluate VLMFP on six domains using GPT-4o (Achiam et al., 2023b). Each problem instance comes with a domain description and an image observation, from which VLMFP generates both PDDL problem and domain files. The problem file is then used as an in-context example to prompt GenVLM to generate PDDL problem files for 100 new instances, with the PDDL Planner validating the plans. We repeat this process for 15 different input instances and calculate the average success rate of finding correct plans, that is, the average success rate of 1500 trials. We follow the same procedure for unseen appearances. Appendix A.5 provides implementation details and failure-case analysis.
Baselines We compare VLMFP against 1) Direct: VLM direct plan generation with an in-context example, 2) CoT: chain-of-thought prompting (Wei et al., 2022) with an in-context example by asking LLMs to reason before generating the final answer, 3) CodePDDL: prompts LLM to generate PDDL files, given problem descriptions generated by SimVLM. For all baselines, we use GPT-4o and also include Direct and CoT baselines with GPT-5 (OpenAI, 2025). All baselines have the same domain description and instance image observation as inputs. For Direct and CoT, since they do not generate PDDL files, we directly apply them to generate plans for the 100 problem instances.
Table 2: Success rate (%) comparison of VLMFP with baselines on 6 grid world domains.
Method Frozenlake Maze Sokoban Package Printer Overcooked Average
S U S U S U S U S U S U S U
[…]
VLMFP $_{\textsc{GPT-4o}}$ 95.2 81.1 88.7 82.8 55.8 25.1 75.2 53.4 58.7 61.0 46.2 21.3 70.0 54.1
Results and Analysis We present VLMFP performance in Table 2. There are three key takeaways:
[…]
Second, VLMFP generalizes well to unseen appearances. On average, VLMFP can solve 54.1% of the unseen problem instances in unseen appearances, which still substantially outperforms the baselines. This demonstrates that the strong perception capability of SimVLM is well adapted by VLMFP, enabling VLMFP to generalize better across different visual styles.
Third, domain complexity influences the relative gains of VLMFP. In simpler domains such as Frozenlake and Maze, both VLMFP and CodePDDL achieve high success rates, though VLMFP still maintains a clear margin. In complex domains like Sokoban and Printer where reasoning about different object types and multiple complicated actions is required, VLMFP delivers a strong performance gain over CodePDDL: 55.8% vs. 0.0% (U) in Sokoban, and 58.7% vs. 7.2% (U) in Printer. This shows that the rating and updating processes are crucial for capturing complex domains.
Together, these findings validate that VLMFP is not only effective and efficient in solving unseen problem instances, but also robust to variations in visual appearances and domain complexity.
4.4 VLMFP Effectiveness of VLMFP Components
We validate each stage of LLMFP with ablations on 6 domains. We examine the effectiveness of Prescreening, Feedback, and Update by removing them from VLMFP one at a time. Similarly, VLMFP generate PDDL files based on one problem instance and test with 100 other problem instances.
Table 3: Success rate (%) when removing key components of VLMFP on 6 grid world domains.
Method Frozenlake Maze Sokoban Package Printer Overcooked Average
No Prescreening 94.3 62.5 23.9 56.9 37.3 10.3 47.5
No Feedback 95.5 84.9 38.0 58.9 53.4 35.9 61.1
No Update 88.1 87.5 0.0 0.0 7.2 1.1 30.7
VLMFP 95.2 88.7 55.8 75.2 58.7 46.2 70.0
We report ablation results in Table 3. Removing any of the three components negatively impacts performance across domains. While prescreening and feedback contribute to steady improvements, the updating stage is essential. Success rates reduce to near zero in complex domains without updating. These demonstrate the effectiveness of all three stages for VLMFP to achieve robust performance.
4.5 Generalization on Game Rules
In previous sections, we evaluate how SimVLM and VLMFP generalize to unseen problem instances and appearances. Beyond these, we also explore whether the framework can adapt to changes in the underlying game rules. In FrozenLake, we created 15 rule variants by modifying core mechanics, such as ice hole behavior or movement dynamics (see Appendix A.6 for all rule descriptions). We fine-tune a SimVLM model with these variations and test it on 5 unseen rules. We test 100 problem instances to examine if SimVLM can provide correct Execution Reason and Result for the steps involving unseen rules: 1) Rule1: Since ice is wet, stepping on an ice hole causes the agent to step forward two steps in the same direction. 2) Rule2: Ice holes are teleports to pos-2-2; 3) Rule3: If you step on an ice hole, you have to execute the same actions three times to actually execute it; 4) Rule4: Stepping on an ice hole unlocks a lucky rocket, where you can step forward three steps in the same direction; 5) Rule5: If you step on an ice hole, you freeze and your next action is skipped.
Table 4: Success rate (%) when testing SimVLM on unseen rules for Frozenlake.
Output Rule1 Rule2 Rule3 Rule4 Rule5
Exec. Reason & Result 94.2 99.0 76.1 59.2 71.1 / 0
The complexity of the five unseen rules increases gradually from Rule1 to Rule5. Rule 1 has the same effects as a training rule (R7 in Appendix A.6), with entirely different wording and explanation. Rule2 changes the teleport destination; the effect is resolved in a single action. Rule3 increases the repeat count from two to three, influencing later actions and requiring multi-step reasoning. Rule4 extends the rocket mechanic to three steps; since the final position is not explicit, it needs extra spatial inference. Results from Rules 1–4 show that SimVLM handles unseen rules in varied formats and altered dynamics with impressive accuracies of 94.2%, 99.0%, 76.1%, and 59.2%.
Rule 5 introduces a completely novel freezing mechanic never observed in training, and here our method fails to generalize. Interestingly, SimVLM correctly states the reasoning as “The agent stepped on ice hole, next action is skipped” but fails to apply it in the next step, yielding 71.1% success in reasoning but 0% in execution. This indicates that while execution on entirely new dynamics remains challenging, SimVLM demonstrates a promising ability to understand and articulate novel rules, suggesting potential to adapt to entirely unseen rules in the future.
5 Conclusion
In this work, we introduced VLMFP, a Dual-VLM-guided framework that autonomously generates PDDL domain and problem files from visual observations. By combining a fine-tuned SimVLM for visual understanding and action simulation with a large GenVLM for symbolic reasoning and iterative refinement, VLMFP removes the need for pre-defined domain files, human supervision, or environment access. Experiments on six grid-world domains show that SimVLM reliably produces scenario descriptions, action simulations, and goal judgments, while VLMFP consistently generates valid PDDL files that enable planners to solve diverse tasks. Results demonstrate strong generalization to unseen instances, appearances, and game rules, underscoring the value of integrating perception and symbolic reasoning for scaling formal planning to complex, real-world domains.
Appendix A Appendix
A.1 Large Language Model Usage for Writing
In this paper, we employed large language models—specifically ChatGPT—solely as writing aids. Draft passages were submitted to the model to improve grammar and refine structure, and the results were carefully reviewed and edited by the authors. The use of the tool was strictly limited to text polishing; it was never applied to generate original content or references.
A.2 Domains Descriptions
We test on six grid world planning domains, Frozenlake, Maze, Sokoban, Package, Printer, and Overcooked (Towers et al., 2024; Silver & Chitnis, 2020; Jin et al., 2023; Wu et al., 2021):
  • Frozenlake In the scenario, you have a girdworld, where each cell can be either normal ground or ice holes. The player starts at a cell, and there is a goal position in a cell. The goal is to move the player to the goal position. You can move up, down, left, right, but you cannot move into the border, and stepping into the ice hole will fail the game.
  • Maze In the scenario, you have a maze gridworld with walls and goal positions. The player can move up, down, left, or right. If the player hits a wall or performs an invalid action, the action fails. The goal is to reach the goal cell.
  • Sokoban In the scenario, you have a gridworld with walls, boxes, and goal positions. The player can move, push the box to goal, and push the box to other position. The player can only push the box forward, not toward other directions. If the player hits a wall or performs an invalid action, the action fails. The goal is to push all boxes to reach the goal cells.
  • Package In the scenario, you have a gridworld with some closed packages in it. The player can turn-left, turn-right, move, pick-up, drop-down, open, or close the packages. If the player hits border or performs an invalid action, the action fails. The goal is to open all packages.
  • Printer In the scenario, you have a gridworld with a desk region and a printer. The player can turn-left, turn-right, move, pick-up, drop-down, toggle-on, or toggle-off the printer. The player cannot move into the desk region. If the player hits border, desk, or performs an invalid action, the action fails. The goal is to pickup and drop the printer in desk region and toggle it on.
  • Overcooked In the scenario, you have a gridworld cooking game with counters, ingredients, chopping boards, and a goal position for delivery. The players can move, chop, pick, drop, merge-ingredient, put-plate, deliver. If the player moves into the counter or performs an invalid action, the action fails. The goal is to cook the salad my merging chopped ingredients, put merged ingredients into plate, and put the plate to the delivery position.
A.3 Scenario Appearance Visualizations
We create six different scenario appearances for Frozenlake, Maze, Sokoban, Package, Printer, and five scenario appearances for Overcooked. We include the visualizations of different appearances in Fig. 4. For Frozenlake, Maze, Sokoban, we change the appearances for every object in the scenario, such as the frozen ice hole, the box, the wall, the agent, etc.. The change objects have various looks. For example, the agent can have the look of a person, a car, or a bike. And the goal can also have the look of a present, a home, a target, or a bag of money. For Package and Printer, we change the color of the background and the desk. For Overcooked, we not only change the color of the background, look of the agents and objects, but also change the ingredients from lettuce and tomato to tomato and onions. The differences in appearances not only bring challenges for the SimVLM to recognize unseen objects, but also to understand the game rule and reason about the objects in the scenario(lettuce or onion?).
Figure 4: Full visualizations of various appearances for 6 grid world domains.
A.4 SimVLM Details
In this section, we will list the dataset statistics, describe the fine-tuning implementation, and show the input and output examples of SimVLM.
A.4.1 Failure case for GPT-5 to accomplish the same task
For the following image observation as input, we show the GPT-5 generated task description and action execution results for the sequence [move left, move down, move down, move up, move up, move right].
Figure 5: The input image to GPT-5 as the image observation of problem instance.
From the output, we can observe that GPT-5 fails to recognize the locations of ice holes. In addition, since it mistakenly thinks pos-2-1 is an ice hole, the action execution also fails. The agent should successfully move to pos-2-1 on Step 2, while GPT-5 marks it as moving into the ice hole, which results in game failure. We include the correct output for this instance from SimVLM in Appendix A.4.4.
From the image we can observe the: Initial State Description:
Initially:
- The size of the gridworld is 4x4.
- The agent is at (pos-1-1).
- The goal is at (pos-4-4).
- The ice holes are at (pos-1-2), (pos-1-3), (pos-2-1), (pos-2-2), (pos-3-2), (pos-3-3).
Your goal is to achieve the following configuration:
- The agent is at (pos-4-4)
Step 1 - Action move left:
Execution Reasoning: The agent tries to move left from pos-1-1. This action is not valid because it would move into the border, so the agent stays at pos-1-1.
Execution result: Unsuccessful.
Step 2 - Action move down:
Execution Reasoning: The agent tries to move down from pos-1-1 to pos-2-1. Cell pos-2-1 is an ice hole, so the agent falls in and the game ends.
Execution result: Unsuccessful.
Step 3 - Action move down:
Execution Reasoning: The agent has already failed the game, so this execution is invalid.
Execution result: Invalid.
Step 4 - Action move up:
Execution Reasoning: The agent has already failed the game, so this execution is invalid.
Execution result: Invalid.
Step 5 - Action move up:
Execution Reasoning: The agent has already failed the game, so this execution is invalid.
Execution result: Invalid.
Step 6 - Action move right:
Execution Reasoning: The agent has already failed the game, so this execution is invalid.
Execution result: Invalid.
Goal reaching: Unsuccessful.
A.4.2 Dataset Statistics
Table 5: Dataset and Task statistics for six grid world domains.
Task FrozenLake Maze Sokoban Package Printer Overcooked
Grid Sizes 3–8 5–8 5–8 4–8 4–8 5–8
#Appearances (S/U) 5/1 5/1 5/1 5/1 5/1 4/1
Action Types 4 4 3 7 7 7
# Data per Appearances 20000 20000 20000 5000 5000 19739
A.4.3 Fine-tuning Implementations
We utilize four H100 80GB GPUs for full-parameter fine-tuning of the Qwen2-VL-7B model. The model is trained for 3 epochs in the SFT stage. The learning rate is set to $1\times 10^{-5}$. We use a batch size of 2 for training. We use 10% of the training data for validation, and we select the model with the lowest validation loss. We perform the inference of SimVLM using a single H100 80GB GPU.
A.4.4 Input and Output Examples
Example Input of the Frozenlake domain:
<image>Given a sequence of actions and an image observation of the initial setup of a scenario, your goal is to decide the execution result after executing given actions. To achieve this task, you will first describe the setup you observe from the image, give the reasoning that explains how the player moves and the success or failure reason for each single action, then rate whether the overall execution is successful, and eventually describe whether the actions achieves the tasks goal.
Task Description: In the scenario, you have a girdworld, where each cell can be either normal ground or ice holes. The left upper corner is (pos-1-1) for (row, column) representation. The player starts at a cell, and there is a goal position in a cell. The goal is to move the player to the goal position. You can move up, down, left, right, but you cannot move into the border, and stepping into the ice hole will fail the game.
Action Sequence:
1: move left
2: move down
3: move down
4: move up
5: move up
6: move right
Example Output of the Frozenlake domain:
From the image we can observe the: Initial State Description:
Initially:
- The size of the gridworld is 4x4.
- The agent is at (pos-1-1).
- The goal is at (pos-4-4).
- The ice holes are at (pos-1-2), (pos-1-4), (pos-2-4), (pos-3-1), (pos-3-3), (pos-3-4), (pos-4-1).
Your goal is to achieve the following configuration:
- The agent is at (pos-4-4)
Step 1 - Action move left:
Execution Reasoning: The agent tries to move left from pos-1-1. This action is not valid as it is moving into the border. So the execution is unsuccessful and the agent stays at the original location.
Execution result: Unsuccessful.
Step 2 - Action move down:
Execution Reasoning: The agent tries to move down from pos-1-1, so it will move to pos-2-1. There is no ice hole in cell pos-2-1 and the agent is not moving to the border, so no invalid action is executed and the agent successfully moves to pos-2-1.
Execution result: Successful.
Step 3 - Action move down:
Execution Reasoning: The agent tries to move down from pos-2-1, so it will move to pos-3-1. Since there is an ice hole in cell pos-3-1, the agent will fall into the ice hole, so the game ended and the execution is unsuccessful.
Execution result: Unsuccessful.
Step 4 - Action move up:
Execution Reasoning: The agent has already failed the game, so this execution is invalid. Execution result: Invalid.
Step 5 - Action move up:
Execution Reasoning: The agent has already failed the game, so this execution is invalid. Execution result: Invalid.
Step 6 - Action move right:
Execution Reasoning: The agent has already failed the game, so this execution is invalid. Execution result: Invalid.
Goal reaching: Unsuccessful
A.5 VLMFP Details
A.5.1 Direct and CoT Major Failure Cases
For all domains, one major failure case for Direct and CoT is because the failure to correctly recognize and locate the objects in the scenario due to insufficient image recognition and spatial reasoning capability. In addition to this, some other failures result from the misunderstanding of actions. For example, in the Package, Printer, and Overcooked domain, when the agent stands in a cell, it can perform actions, such as open and close, to objects in adjacent cells. But the VLMs often think the agent should stand in the same cell as the object to operate it. Including in-context examples is beneficial, but could not stop all failures like this.
A.5.2 VLMFP Major Failure Cases
For failure in generating correct PDDL problem files, one major failure case for all domains is that it fails to generate all states that describe directional relationships between cells. For example, ‘move-dir pos-1-1 pos-1-2 right’ defines ‘pos-1-2 is at right to pos-1-1’. Although the problem description is correct, the GenVLM sometimes fails to include all available directional relationships, so that some actions cannot be performed. In addition to this type of failure, for complex domains, there are often different types of objects. Sometimes, GenVLM fails to assign types for all objects.
For failure in generating correct Domain problem files, the major failure case is failure to describe correct action preconditions and effects. For example, in the Package domain, the agent can only open the package it faces. The definition for the open action is:
(:action open
:parameters (?pkg - package ?pos - position ?pkgpos - position ?dir - direction)
:precondition (and
(at ?pos)
(package-at ?pkg ?pkgpos)
(package-closed ?pkg)
(facing ?dir)
(move-dir ?pos ?pkgpos ?dir)
:effect (and
(not (package-closed ?pkg))
(package-open ?pkg)
This ‘facing ?dir’ precondition and the ‘(move-dir ?pos ?pkgpos ?dir)’ are easy to neglect.
For CodePDDL, it has similar failure cases as VLMFP. Without the updating process, the failures cannot be corrected.
A.5.3 VLMFP Implementation Details and Prompts
Prescreening During prescreening, GenVLM first generates the problem PDDL file based on the domain description, problem description, and image observation. The descriptions include the problem and domain PDDL template, which only includes the available object types, action names, and action variable names. In the prompt, a non-task-specific example is included. The example is a 3x3 grid world with no obstacles. The prompt is:
Given a natural language description of a planning problem and an image observation of the initial scenario, your task is to generate a complete PDDL problem instance that is equivalent to its natural language description and image observation, contain all necessary predicates as described in the description, and is thorough and complete. Consider all predicates that are relevant and neccesary.
Example:
Domain Description:
{domain_nl_wrapped}
Domain PDDL Template:
‘‘‘pddl
(define (domain grid)
(:requirements :strips)
(:predicates)
(:action move-up
:parameters (?from ?to)
:precondition ()
:effect ()
(:action move-down
:parameters (?from ?to)
:precondition ()
:effect ()
(:action move-left
:parameters (?from ?to)
:precondition ()
:effect ()
(:action move-right
:parameters (?from ?to)
:precondition ()
:effect ()
Problem Description:
‘‘‘markdown
You are tasked with manipulating an agent to reach the goal without colliding into the wall. The position representation is (row, column) representation. For example, (pos-4-3) represents the position in fourth row and third column. The left upper corner is (pos-1-1). You can perform four actions: move-up, move-down, move-left, and move-right.
Initially:
- The agent is at (pos-2-1).
- The goal is at (pos-3-3).
Your goal is to achieve the following configuration:
- The agent is at (pos-3-3)
Problem PDDL Template:
‘‘‘pddl
(define (problem grid)
(:domain grid)
(:objects )
(:init )
(:goal (and ))
Problem PDDL:
‘‘‘markdown
(define (problem grid) (:domain grid)
(:objects
pos-1-1 - position
pos-1-2 - position
pos-1-3 - position
pos-2-1 - position
pos-2-2 - position
pos-2-3 - position
pos-3-1 - position
pos-3-2 - position
pos-3-3 - position
(:init
(at pos-2-1)
(is-goal pos-3-3)
(move-dir-left pos-1-2 pos-1-1)
(move-dir-left pos-1-3 pos-1-2)
(move-dir-left pos-2-2 pos-2-1)
(move-dir-left pos-2-3 pos-2-2)
(move-dir-left pos-3-2 pos-3-1)
(move-dir-left pos-3-3 pos-3-2)
(move-dir-right pos-1-1 pos-1-2)
(move-dir-right pos-1-2 pos-1-3)
(move-dir-right pos-2-1 pos-2-2)
(move-dir-right pos-2-2 pos-2-3)
(move-dir-right pos-3-1 pos-3-2)
(move-dir-right pos-3-2 pos-3-3)
(move-dir-up pos-2-1 pos-1-1)
(move-dir-up pos-2-2 pos-1-2)
(move-dir-up pos-2-3 pos-1-3)
(move-dir-up pos-3-1 pos-2-1)
(move-dir-up pos-3-2 pos-2-2)
(move-dir-up pos-3-3 pos-2-3)
(move-dir-down pos-1-1 pos-2-1)
(move-dir-down pos-1-2 pos-2-2)
(move-dir-down pos-1-3 pos-2-3)
(move-dir-down pos-2-1 pos-3-1)
(move-dir-down pos-2-2 pos-3-2)
(move-dir-down pos-2-3 pos-3-3)
(:goal (at pos-3-2))
Domain Description:
{target_domain_nl}
Domain PDDL Template:
{target_domain_template_pddl}
Problem Description:
{target_problem_nl}
Problem PDDL Template:
{target_problem_template_pddl}
After generating the problem file, GenVLM continues to generate the domain file based on the example of the grid, domain, and problem descriptions, image observation, and the generated problem file. The prompt for this is:
You are given a natural language description of a planning problem in the domain {target_domain_name} along with one problem instance in PDDL format and an image observation of the initial setup. Your task is to generate a PDDL domain for the target domain {target_domain_name} that is equivalent to its natural language description and is compatible with the provided problem instance.
Starting from a PDDL domain template, you are allowed to modify the template using the following two python function interfaces:
‘‘‘python
add_or_update_predicates(predicates: List[str])
modify_action(action_name: str, new_preconditions: List[str], new_effects: List[str])
An example of above functions applied to an example PDDL domain template is as follows:
Example Domain Description:
‘‘‘markdown
The robot has four actions: move-up, move-down, move-left, and move-right. This domain models an agent navigating in a grid world. The goal is to reach a target location. The world consists of a set of discrete positions (e.g., grid cells). The position representation is (row, column) representation. For example, (pos-4-3) represents the position in fourth row and third column, and (pos-4-4) is at right of (pos-4-3). Between each two positions, the directional links between them design how they can move between each other. For example, the agent can move-left from pos-3-2 to pos-3-1 because pos-3-1 is at left to pos-3-2, but cannot move-left from pos-3-2 to pos-3-3.
Example Problem PDDL:
‘‘‘pddl
(define (problem maze) (:domain maze)
(:objects
pos-1-1 - position
pos-1-2 - position
pos-1-3 - position
pos-2-1 - position
pos-2-2 - position
pos-2-3 - position
pos-3-1 - position
pos-3-2 - position
pos-3-3 - position
(:init
(at pos-2-1)
(is-goal pos-3-3)
(move-dir-left pos-1-2 pos-1-1)
(move-dir-left pos-1-3 pos-1-2)
(move-dir-left pos-2-2 pos-2-1)
(move-dir-left pos-2-3 pos-2-2)
(move-dir-left pos-3-2 pos-3-1)
(move-dir-left pos-3-3 pos-3-2)
(move-dir-right pos-1-1 pos-1-2)
(move-dir-right pos-1-2 pos-1-3)
(move-dir-right pos-2-1 pos-2-2)
(move-dir-right pos-2-2 pos-2-3)
(move-dir-right pos-3-1 pos-3-2)
(move-dir-right pos-3-2 pos-3-3)
(move-dir-up pos-2-1 pos-1-1)
(move-dir-up pos-2-2 pos-1-2)
(move-dir-up pos-2-3 pos-1-3)
(move-dir-up pos-3-1 pos-2-1)
(move-dir-up pos-3-2 pos-2-2)
(move-dir-up pos-3-3 pos-2-3)
(move-dir-down pos-1-1 pos-2-1)
(move-dir-down pos-1-2 pos-2-2)
(move-dir-down pos-1-3 pos-2-3)
(move-dir-down pos-2-1 pos-3-1)
(move-dir-down pos-2-2 pos-3-2)
(move-dir-down pos-2-3 pos-3-3)
(:goal (at pos-3-2))
Example PDDL Template:
‘‘‘pddl
(define (problem maze)
(:domain maze)
(:objects )
(:init )
(:goal (and ))
Example Completion:
‘‘‘python
add_or_update_predicates([
’(move-dir-up ?x ?y)’,
’(move-dir-down ?x ?y)’,
’(move-dir-left ?x ?y)’,
’(move-dir-right ?x ?y)’,
’(at ?x)’,
’(is-goal ?x)’
modify_action(’move-up’, [
"(at ?from)",
"(move-dir-up ?from ?to)"
"(not (at ?from))",
"(not (clear ?to))",
"(at ?to)",
modify_action(’move-down’, [
"(at ?from)",
"(move-dir-down ?from ?to)"
"(not (at ?from))",
"(at ?to)",
modify_action(’move-left’, [
"(at ?from)",
"(move-dir-left ?from ?to)"
"(not (at ?from))",
"(at ?to)",
modify_action(’move-right’, [
"(at ?from)",
"(move-dir-right ?from ?to)"
"(not (at ?from))",
"(at ?to)",
Target Domain Description:
{target_domain_nl}
Target Problem PDDL:
{target_problem_pddl}
Now, your task is to complete the following PDDL template by generating necessary predicates and action preconditions and effects:
Target PDDL Template:
{target_domain_template_pddl}
You must never modify action parameters, and you are only allowed to use the following two function interfaces to modify the template.
PDDL files updating During updating, GenVLM takes in the descriptions, image observation, previously generated problem and domain PDDL files, and execution inconsistency feedback. GenVLM is prompted to reason about which part of the files is problematic and then update the file(s).
A.6 Genelization to Game Rules
A.6.1 Game Rule Variation Descriptions
We designed 15 new rules and collected data for these rules under the default Frozenlake appearance, with size 3-8, and different ice hole probabilities. For each rule, we collect 20k data. Here’s the detailed rule descriptions for the new rules.
  • R1: Ice holes do not result in failure. Instead, the agent must step on one ice hole once as a precondition for game completion.
  • R2: Variant of R1. The agent must step on two ice holes as a precondition for game completion.
  • R3: Ice holes function as teleports. There are two ice holes in the scenario that allow teleportation between each other.
  • R4: Ice holes function as teleports. Any ice hole allows the agent to teleport back to the origin position.
  • R5: If the agent steps on an ice hole, it must execute the same action twice to actually execute it.
  • R6: The agent has two lives. It does not fail when stepping on the first ice hole, but fails the game when stepping on the second one.
  • R7: Stepping on an ice hole unlocks a lucky rocket, allowing the agent to step forward two steps in the same direction.
  • R8: Ice is slippery. If the agent steps on ice and the next cell in the same direction is also ice, the agent slips to the second ice position, continuing to slip until reaching a non-ice position.
  • R9: The agent can only step into an ice hole if entering from above. Stepping in from other directions is invalid.
  • R10: After stepping on an ice hole, the agent must always execute actions twice to actually execute them.
  • R11: Variant of R10. The agent must execute actions three times instead of twice.
  • R12: If the agent steps onto ice, it slides in that direction until hitting a wall.
  • R13: Variant of R12. The agent bounces back until reaching a wall instead of sliding forward.
  • R14: Stepping on an ice hole swaps the goal and origin positions.
  • R15: The agent can only move up or down when on ice holes.
A.7 Ground truth PDDL files examples
Here we show ground truth PDDL Problem and Domain file examples for the Frozenlake and Package domain. They represent very different sets of actions in grid world domains. In Frozenlake, the agent does not have orientations; that is, it can freely move to any adjacent cell. However, in the Package domain, the agent has orientation and can only move in the direction it faces. Thus, the move action definition is different, and the turn action is also required in the Package domain. In addition, the Package domain also shows some manipulation actions, which are common in complex grid world domains.
Frozenlake Sample PDDL Problem File:
(define (problem FL-rand)
(:domain frozenlake)
(:objects pos-1-1 pos-1-2 pos-1-3 pos-1-4 pos-2-1 pos-2-2 pos-2-3 pos-2-4 pos-3-1 pos-3-2 pos-3-3 pos-3-4 pos-4-1 pos-4-2 pos-4-3 pos-4-4)
(:init
(at pos-1-1)
(ice-hole pos-1-3)
(ice-hole pos-1-4)
(ice-hole pos-2-2)
(ice-hole pos-3-3)
(ice-hole pos-3-4)
(ice-hole pos-4-1)
(up_direction pos-2-1 pos-1-1)
(up_direction pos-2-2 pos-1-2)
(up_direction pos-2-3 pos-1-3)
(up_direction pos-2-4 pos-1-4)
(up_direction pos-3-1 pos-2-1)
(up_direction pos-3-2 pos-2-2)
(up_direction pos-3-3 pos-2-3)
(up_direction pos-3-4 pos-2-4)
(up_direction pos-4-1 pos-3-1)
(up_direction pos-4-2 pos-3-2)
(up_direction pos-4-3 pos-3-3)
(up_direction pos-4-4 pos-3-4)
(down_direction pos-1-1 pos-2-1)
(down_direction pos-1-2 pos-2-2)
(down_direction pos-1-3 pos-2-3)
(down_direction pos-1-4 pos-2-4)
(down_direction pos-2-1 pos-3-1)
(down_direction pos-2-2 pos-3-2)
(down_direction pos-2-3 pos-3-3)
(down_direction pos-2-4 pos-3-4)
(down_direction pos-3-1 pos-4-1)
(down_direction pos-3-2 pos-4-2)
(down_direction pos-3-3 pos-4-3)
(down_direction pos-3-4 pos-4-4)
(left_direction pos-1-2 pos-1-1)
(left_direction pos-1-3 pos-1-2)
(left_direction pos-1-4 pos-1-3)
(left_direction pos-2-2 pos-2-1)
(left_direction pos-2-3 pos-2-2)
(left_direction pos-2-4 pos-2-3)
(left_direction pos-3-2 pos-3-1)
(left_direction pos-3-3 pos-3-2)
(left_direction pos-3-4 pos-3-3)
(left_direction pos-4-2 pos-4-1)
(left_direction pos-4-3 pos-4-2)
(left_direction pos-4-4 pos-4-3)
(right_direction pos-1-1 pos-1-2)
(right_direction pos-1-2 pos-1-3)
(right_direction pos-1-3 pos-1-4)
(right_direction pos-2-1 pos-2-2)
(right_direction pos-2-2 pos-2-3)
(right_direction pos-2-3 pos-2-4)
(right_direction pos-3-1 pos-3-2)
(right_direction pos-3-2 pos-3-3)
(right_direction pos-3-3 pos-3-4)
(right_direction pos-4-1 pos-4-2)
(right_direction pos-4-2 pos-4-3)
(right_direction pos-4-3 pos-4-4)
(:goal
(and
(at pos-4-4)
Frozenlake PDDL Domain File:
(define (domain frozenlake)
(:requirements :strips)
(:predicates (at ?x) (down_direction ?from ?to) (ice-hole ?x) (left_direction ?from ?to) (right_direction ?from ?to) (up_direction ?from ?to))
(:action move-down
:parameters (?from ?to)
:precondition (and (at ?from) (down_direction ?from ?to) (not (ice-hole ?from)))
:effect (and (at ?to) (not (at ?from)))
(:action move-left
:parameters (?from ?to)
:precondition (and (at ?from) (left_direction ?from ?to) (not (ice-hole ?from)))
:effect (and (at ?to) (not (at ?from)))
(:action move-right
:parameters (?from ?to)
:precondition (and (at ?from) (right_direction ?from ?to) (not (ice-hole ?from)))
:effect (and (at ?to) (not (at ?from)))
(:action move-up
:parameters (?from ?to)
:precondition (and (at ?from) (up_direction ?from ?to) (not (ice-hole ?from)))
:effect (and (at ?to) (not (at ?from)))
Package Sample PDDL Problem File:
(define (problem package)
(:domain package)
(:objects
; Positions in 4x4 grid
pos-1-1 pos-1-2 pos-1-3 pos-1-4 pos-2-1 pos-2-2 pos-2-3 pos-2-4 pos-3-1 pos-3-2 pos-3-3 pos-3-4 pos-4-1 pos-4-2 pos-4-3 pos-4-4 - position
; Packages
pkg-1 pkg-2 - package
; Directions
up down left right - direction
(:init
; Agent initial position and orientation
(at pos-3-3)
(facing up)
; Package locations and states
(package-at pkg-1 pos-1-3)
(package-closed pkg-1)
(package-at pkg-2 pos-4-1)
(package-closed pkg-2)
; Turn relations
(left-turn up left)
(left-turn left down)
(left-turn down right)
(left-turn right up)
(right-turn up right)
(right-turn right down)
(right-turn down left)
(right-turn left up)
; Grid adjacency relationships
(move-dir pos-1-1 pos-1-2 right) (move-dir pos-1-2 pos-1-1 left)
(move-dir pos-1-2 pos-1-3 right) (move-dir pos-1-3 pos-1-2 left)
(move-dir pos-1-3 pos-1-4 right) (move-dir pos-1-4 pos-1-3 left)
(move-dir pos-2-1 pos-2-2 right) (move-dir pos-2-2 pos-2-1 left)
(move-dir pos-2-2 pos-2-3 right) (move-dir pos-2-3 pos-2-2 left)
(move-dir pos-2-3 pos-2-4 right) (move-dir pos-2-4 pos-2-3 left)
(move-dir pos-3-1 pos-3-2 right) (move-dir pos-3-2 pos-3-1 left)
(move-dir pos-3-2 pos-3-3 right) (move-dir pos-3-3 pos-3-2 left)
(move-dir pos-3-3 pos-3-4 right) (move-dir pos-3-4 pos-3-3 left)
(move-dir pos-4-1 pos-4-2 right) (move-dir pos-4-2 pos-4-1 left)
(move-dir pos-4-2 pos-4-3 right) (move-dir pos-4-3 pos-4-2 left)
(move-dir pos-4-3 pos-4-4 right) (move-dir pos-4-4 pos-4-3 left)
(move-dir pos-1-1 pos-2-1 down) (move-dir pos-2-1 pos-1-1 up)
(move-dir pos-1-2 pos-2-2 down) (move-dir pos-2-2 pos-1-2 up)
(move-dir pos-1-3 pos-2-3 down) (move-dir pos-2-3 pos-1-3 up)
(move-dir pos-1-4 pos-2-4 down) (move-dir pos-2-4 pos-1-4 up)
(move-dir pos-2-1 pos-3-1 down) (move-dir pos-3-1 pos-2-1 up)
(move-dir pos-2-2 pos-3-2 down) (move-dir pos-3-2 pos-2-2 up)
(move-dir pos-2-3 pos-3-3 down) (move-dir pos-3-3 pos-2-3 up)
(move-dir pos-2-4 pos-3-4 down) (move-dir pos-3-4 pos-2-4 up)
(move-dir pos-3-1 pos-4-1 down) (move-dir pos-4-1 pos-3-1 up)
(move-dir pos-3-2 pos-4-2 down) (move-dir pos-4-2 pos-3-2 up)
(move-dir pos-3-3 pos-4-3 down) (move-dir pos-4-3 pos-3-3 up)
(move-dir pos-3-4 pos-4-4 down) (move-dir pos-4-4 pos-3-4 up)
(:goal
(and
(package-open pkg-1)
(package-open pkg-2)
Package PDDL Domain File:
(define (domain package)
(:requirements :strips :typing)
(:types
position package direction
(:predicates
(at ?pos - position) ; agent is at position
(package-at ?pkg - package ?pos - position) ; package is at position
(package-open ?pkg - package) ; package is open
(package-closed ?pkg - package) ; package is closed
(facing ?dir - direction) ; agent is facing direction
(left-turn ?from - direction ?to - direction)
(right-turn ?from - direction ?to - direction)
(move-dir ?pos1 - position ?pos2 - position ?dir - direction) ; positions are move-dir in direction
(:action turn-left
:parameters (?current-dir - direction ?new-dir - direction)
:precondition (and
(facing ?current-dir)
(left-turn ?current-dir ?new-dir)
:effect (and
(not (facing ?current-dir))
(facing ?new-dir)
(:action turn-right
:parameters (?current-dir - direction ?new-dir - direction)
:precondition (and
(facing ?current-dir)
(right-turn ?current-dir ?new-dir)
:effect (and
(not (facing ?current-dir))
(facing ?new-dir)
(:action move
:parameters (?from - position ?to - position ?dir - direction)
:precondition (and
(at ?from)
(facing ?dir)
(move-dir ?from ?to ?dir)
:effect (and
(not (at ?from))
(at ?to)
(:action open
:parameters (?pkg - package ?pos - position ?pkgpos - position ?dir - direction)
:precondition (and
(at ?pos)
(package-at ?pkg ?pkgpos)
(package-closed ?pkg)
(facing ?dir)
(move-dir ?pos ?pkgpos ?dir)
:effect (and
(not (package-closed ?pkg))
(package-open ?pkg)
(:action close
:parameters (?pkg - package ?pos - position ?pkgpos - position ?dir - direction)
:precondition (and
(at ?pos)
(package-at ?pkg ?pkgpos)
(package-open ?pkg)
(facing ?dir)
(move-dir ?pos ?pkgpos ?dir)
:effect (and
(not (package-open ?pkg))
(package-closed ?pkg)